OpenStudy (anonymous):

Linear Equations in One Variable Translate Word Problems into System of Equations Eight boys rented a motor boat for their outing to puerto galera . three of them backed out at the last minute. As a result each of the remaining five boys had to pay off 90$ more that they had expected to pay how much was the rental of the boat?

OpenStudy (anonymous):

Let x be the portion of the rent for each boy and rent be the rent. Solve the following for x and rent. \[\{8x=\text{rent},5(x+90)=\text{rent}\} \]{rent = 1200, x = 150}

OpenStudy (anonymous):

thanks rob, but how did you get it?

OpenStudy (anonymous):

im trying to find out how to solve this because i miss this lesson earlier.

OpenStudy (anonymous):

Are you unsure how to set up the equations or solve them? Thank you for the medal by the way.

OpenStudy (anonymous):

8x=5(x+90) Can you solve the above for x?

OpenStudy (anonymous):

8x=5x-450 x-5x=450-8? im not sure.

OpenStudy (anonymous):

8x=5(x+90) 8x=5x+450 3x=450 x=150

OpenStudy (anonymous):

o i get it now.

OpenStudy (anonymous):

how about the 1200?

OpenStudy (anonymous):

150x8?

OpenStudy (anonymous):

8x=rent 8*150=rent 1200=rent

OpenStudy (anonymous):

That all there is. Actually, I love word problems.

OpenStudy (anonymous):

ok ill pm you if i need help thanks rob.

OpenStudy (anonymous):

Have a good day.

OpenStudy (anonymous):

yes you too

OpenStudy (anonymous):

rob may i ask last question how did 3 become x? 3x=450 x=150

OpenStudy (amistre64):

there are some fundamental properties called identity and inverse identity is such that a*id = a , id would have to be 1 inverse is such that a*in = 1 , in would have to be 1/a, given that a not equal 0

OpenStudy (amistre64):

3x = 450 , the inverse of 3 is 1/3 /3 /3 ---------- 1x = 150 , 1x = x x = 150

OpenStudy (anonymous):

oh. thanks now i get it.

OpenStudy (amistre64):

good luck ;)

OpenStudy (anonymous):

thank you